OpenStudy (anonymous):

What is the product? 2/3 x 5/8 x 3/4 A. 5/16 B. 5/8 C. 7/12 D. 3/4

OpenStudy (dariusx):

2x5x3 / 3x8x4 = 30 / 96 = 15 / 48 = 5 / 16

OpenStudy (kendricklamar2014):

\[\frac{ 2 }{ 3 } + \frac{ 5 }{ 8 } = \frac{ 5 }{ 12 }\] \[\frac{ 5 }{ 12 } + \frac{ 3 }{ 4 } = \frac{ 5 }{ 16 }\]

OpenStudy (vickyisthesmartone123):

She's correct^